x+4y=-7
-4x+5y=28

Answer by dragonwalker(73)

okay, so let us call x + 4y = -7 EQUATION ONE
and -4x + 5y = 28 EQUATION TWO
from equation one you can get:
x = -7 -4y (by moving the 4y to the other side (and when doing so changing the sign).
You can then multiply all values by four to give:
4x = -28 - 16y
Now change the signs for all values to get -4x:
-4x = 28 + 16y
you can then substitute this into equation 2:
(28 + 16y) + 5y = 28
16y + 5y = 28 - 28
21y = 0
so y = 0
Then use equation one with a value of y=0:
x = -7
so x= -7 and y=0